# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v4

@@ 3.2 Identification of the shocks by non-Gaussianity
    The structural error process $e_{t}=(e_{1t},...,e_{dt})$ is a sequence of independent and identically distributed random vectors such that $e_{t}$ is independent of $\mathcal{F}_{t-1}$ and each component $e_{it}$, $i=1,...,d$, follows a continuous distribution with zero mean and unit variance.
@@
    The components of $e_{t}=(e_{1t},...,e_{dt})$ are mutually independent and at most one them has a Gaussian marginal distribution.
Note that the assumption of a unit variance is merely a normalization since the variance of the shocks is captured by the impact matrix $B_{y,t}$. Also, due to the time-variation of the impact matrix (3.3), Assumption 1 does not imply that the reduced form innovations $u_{t}$ are independent nor that they are identically distributed.
The next lemma then shows that at each $t$, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is uniquely identified up to ordering and signs of its columns. This lemma, which is partly similar to Proposition 1 in Lanne et al. (2017), is used in the following sections to show that the impact matrices $B_{1},...,B_{M}$ of the regimes in (3.3) are identified. Moreover, Lemma 2 also shows that if the transition weights are either exogenous and nonrandom, or they follow the threshold (2.3) or the logistic (2.4) process, then also the autoregressive (AR) and weight function parameters are identified. Specific weight functions are assumed here for simplicity, but the result can be extended to other suitable weight functions as well.
@@
Consider the STVAR model defined in Equations (3.2) and (3.3) with Assumption 1 satisfied. Then, conditionally on $\mathcal{F}_{t-1}$, at each $t$, $B_{y,t}$ is uniquely identified up to ordering and signs of its columns almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
Moreover, suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$ and that the following conditions hold:
  1. (1)
    the transition weights $\alpha_{m,t}$, $m=1,...,M$, are either exogenous (nonrandom), or they follow the threshold (2.3) or the logistic process (2.4) (with $M=2$ in the latter),
@@
    there exists $M$ indices $t_{1},...,t_{M}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, $i=1,...,M$, are linearly independent.
Then, for models with exogenous weights, the parameters $\phi_{m},A_{m,1},...A_{m,p}$, $m=1,...,M$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$. For threshold models (2.3), the parameters $\phi_{m},A_{m,1},...A_{m,p}$, $m=1,...,M$ and $r_{1},...,r_{m-1}$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$. For logistic models (2.4), the parameters $\phi_{m},A_{m,1},...A_{m,p}$, $m=1,2$, and $c,\gamma$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ defined in Section 2 is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
Lemma 2 is proven in Appendix B.2.
We have concluded that at each $t$, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is unique up to ordering and signs of its columns under Assumption 1. In other words, at each $t$‚ changing the ordering or signs of the columns of $B_{y,t}$ would lead to an observationally equivalent model, but changing $B_{y,t}$ in any other way would lead to an observationally distinct model. The result holds only almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, because $B_{y,t}$ is invertible almost everywhere by Lemma 1. Identification of the AR and weight function parameters, in turn, is achieved only almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ (and $(z_{1},...,z_{T})\in\mathbb{R}^{T}$), because the identification fails in some special cases that comprise a set with Lebesgue measure zero. Condition (3) of Lemma 2 ensures that the transition weights exhibit enough variation to distinguish between different parameter values. If there are only two regimes, this condition is satisfied if transition weights take at least two distinct values.
If the impact matrix is time-invariant as in Lanne et al. (2017), i.e., $B_{y,t}=B$ for some constant matrix $B$, it follows from Lemma 2 that the structural shocks are identified up to ordering and signs. Such statistically identified shocks can then often be labelled as economic shocks based on external information (see Lanne et al., 2017). However, when the impact matrix varies in time, two complications arise. First, as the impact matrix $B_{y,t}$ is not a matrix of constant parameters but a function of parameters, it needs to be shown that the parameters in the specific functional form of $B_{y,t}$ are identified. Second, since the identification of $B_{y,t}$ at each $t$ is only up to ordering and signs of its columns, its unique identification requires that the parameters in the functional form of $B_{y,t}$ are restricted to fix the ordering and signs of the columns of $B_{y,t}$.
To address these complications, consider our functional form of $B_{y,t}$ in Equation (3.3), which defines $B_{y,t}$ as a weighted sum of the impact matrices of the regimes $B_{1},...,B_{M}$ with the weights given by the transition weights $\alpha_{m,t}$, $m=1,...,M$. To fix the ordering and signs of the columns of $B_{y,t}$, such constraints need to be imposed on $B_{1},...,B_{M}$ that under them, the impact matrices $B_{y,t}$ before and after reordering or changing the signs of the columns of any of $B_{1},...,B_{M}$ are observationally distinct at some $t$. However, constraints that are not restrictive in some specifications may be overidentifying in others, while constraints sufficient in some specifications may not be identifying in others, depending on the transition weights and shock distributions. Therefore, it is useful to divide the discussion on identification into the cases of binary (i.e., $\alpha_{m,t}\in\{0,1\}$) and more general transition weights.
3.3 Identification of the shocks in structural TVAR models
Suppose the transition weights are binary, $\alpha_{m,t}\in\{0,1\}$, for all $t$ and $m=1,...,M$, specifically being either exogenous or of the threshold form (2.3). Then at every $t$, the process is completely in one of the regimes, making the impact matrix $B_{y,t}=B_{m}$ for the active regime $m$ with $\alpha_{m,t}=1$. By Lemma 2, $B_{y,t}$ is thereby identified up to ordering and signs of its columns within each regime, and identification requires constraints that fix these orderings and signs for each regime-specific impact matrix $B_{1},...,B_{M}$. For simplicity, and to state the identification result under just-identifying constraints, we assume each shock $e_{it}$, $i=1,...,d$, follows the skewed $t$-distribution of Hansen (1994), which can flexibly capture fat tails and skewness, and nests the Gaussian distribution as a special case. The skewed $t$-distribution is described in Section 4 discussing estimation of the model parameters. [^1] The identification result can be extended to other suitable shock distributions using arguments similar to those in the proof of Proposition 1 in Appendix B.3.
The following proposition, which is proved in Appendix B.3, states the identification result for the TVAR model and models with exogenous binary weights.
@@ 3.3 Identification of the shocks in structural TVAR models
Consider the STVAR model defined in Equations (3.2) and (3.3) with $B_{1},...,B_{M}$ invertible; Assumption 1 and the conditions of Lemma 2 satisfied; $\alpha_{m,t}$ either of the threshold form (2.3) or exogenous (nonrandom) with $\alpha_{m,t}\in\{0,1\}$ for all $t$ and $m=1,...,M$; and the shocks $e_{it}$, $i=1,...,d$, following the skewed $t$-distribution defined in Equation (4.1) with the parameters $\nu_{i}>2$ and $\lambda_{i}\in(-1,1)$ controlling kurtosis and skewness, respectively. Suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, and assume the following conditions hold:
  1. (a)
    the ordering and signs of the columns of $B_{1}$ are fixed,
  2. (b)
    $(\nu_{i},\lambda_{i})\neq(\nu_{j},\lambda_{j})$ for all $i\neq j\in\{1,...,d\}$,
  3. (c)
    $\lambda_{i}\neq 0$ for all $i=1,...,d$.
Then, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T})\in\mathbb{R}^{d(T+p)}$.
Condition (a) of Proposition 1 fixes the ordering and signs of $B_{1}$, which effectively fixes the ordering and signs of the columns of $B_{2},...,B_{M}$ under Conditions (b) and (c), respectively. Any fixed ordering and signs of the columns of $B_{1}$ can be chosen without loss of generality (e.g., by making the first nonzero entry in each column is positive and ordering these entries in decreasing magnitude, given that none of them are equal). If Condition (b) fails for some $i,j\in\{1,...,d\}$, the distributions of these shocks are identical, implying that fixing the ordering of the columns of $B_{1}$ does not fix the ordering of the corresponding columns of $B_{2},...,B_{M}$. In such a case, the ordering of the corresponding columns of $B_{2},...,B_{M}$ can be arbitrarily fixed without loss of generality. Similarly, if Condition (c) fails for some $i\in\{1,...,d\}$ (meaning zero skewness), the signs of the corresponding columns of $B_{2},...,B_{M}$ can be arbitrarily fixed without loss of generality. The result of Proposition 1 holds only almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T})\in\mathbb{R}^{d(T+p)}$, because in some special cases that comprise set with a Lebesgue measure zero, changing the ordering of the columns of $B_{2},...,B_{M}$ does not change the joint density of the model.
If Condition (b) or (c) is close to break, identification is weak with respect to the ordering or signs of the corresponding columns of $B_{2},...,B_{M}$, respectively. Somewhat surprisingly, our experience shows that even when Conditions (b) and (c) hold clearly, identification appears to be often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, at least if the parameters are estimated by the method of maximum likelihood or penalized maximum likelihood (discussed in Section 4). Specifically, there are frequently a number of local maxima of the (penalized) log-likelihood function with (penalized) log-likelihoods close to each other, corresponding to otherwise largely similar parameter estimates but with different the ordering and signs of the columns of $B_{2},...,B_{M}$. As discussed in Section 3.5, this weak identification should be appropriately dealt with when labelling the shocks, for instance, by combining Proposition 1 with additional identifying information.
3.4 Identification of the shocks in structural STVAR models
Having established the identification result for structural TVAR models, we now consider transition weights that exhibit more variation. In particular, the transition weights are here assumed to be either exogenous (nonrandom) or logistic as defined in (2.4), so that the result of Lemma 2 establishing identification of the transition weights can be exploited. Under the conditions of the following proposition (which is proven in Appendix B.4), the matrices $B_{1},...,B_{M}$ are uniquely identified.
Proposition 2.
Consider the STVAR model defined in Equations (3.2) and (3.3) with Assumption 1 and the conditions of Lemma 2 satisfied, and let the transition weights be either exogenous (nonrandom) or follow the logistic process defined in (2.4) (with $M=2$ assumed). Suppose the observed time series is $y_{-p+1},...,y_{0},y_{1},...,y_{T}$, and assume the following conditions hold:
@@ 3.4 Identification of the shocks in structural STVAR models
    the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero.
@@
Then, for models with logistic transition weights, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$. For models with nonrandom exogenous transition weights, this result holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
@@
Proposition 2 essentially states that if the structural shocks are independent and at most one of them is Gaussian (Assumption 1), the impact matrices $B_{1},...,B_{M}$, and hence, the structural shocks are identified. The result holds only almost everywhere, because the parameters are not identified in some special cases that constitute a set with Lebesgue measure zero (see also Lemma 2).
@@
Condition (A) of Proposition 2 states that the ordering and signs of the columns of $B_{1}$ should be fixed, which fixes the ordering and signs of the columns of $B_{y,t}$. If the shocks follow the skewed $t$-distributions defined in (4.1), any fixed ordering and signs of the columns of $B_{1}$ can be assumed without loss of generality, as reordering the columns would just reorder the shocks and swapping a sign of a column corresponds to swapping the sign of the related shock and skewness parameter. Condition (B) states that the transition weights should, for some $t$, be strictly positive for all of the regimes. If there are only two regimes, the requirement is merely that the regime-switches are not always discrete, in which case Proposition 1 would be relevant. Similarly to the structural TVAR model discussed in Section 3.3, our experience shows that identification appears to be often weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, which should be appropriately dealt with when labelling the shocks, as discussed in the next section.
3.5 Labelling the shocks and blended identification
As in the linear SVAR model of Lanne et al. (2017), the statistically identified structural shocks do not necessarily have an economic interpretation, and labelling them as economic shocks requires external information. Our identification method, nonetheless, facilitates identifying structural shocks with time-varying impact effects without imposing restrictions on the impact responses of the variables. This is a major advantage compared to statistical identification by heteroscedasticity that restricts how the impact responses may vary over time (see Lütkepohl and Netšunajev, 2017).
Labelling the shocks based on the estimates of the impact matrices $B_{1},...,B_{M}$ might not always be straightforward, as the same shock needs to be associated with the same column of the impact matrix $B_{m}$ in all regimes. For example, if a positive supply shock increases output and decreases prices on impact, and it should do so in all regimes, labelling the $i$th shock as the supply shock requires the $i$th column of all $B_{1},...,B_{M}$ to satisfy such signs. However, as discussed in Sections 3.3 and 3.4, identification can be weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, which should be appropriately accounted for when labelling the shocks.
Assuming that the parameters are estimated by the method of maximum likelihood (ML) or penalized maximum likelihood (PML) described in Section 4, weak identification often results in multiple local maxima of the (penalized) log-likelihood function with the related (penalized) log-likelihoods close to each other. If identification is weak with respect to the ordering and signs of the columns of $B_{2},...,B_{M}$, these local solutions often correspond to otherwise largely similar parameter estimates but with different ordering and signs of the columns of $B_{2},...,B_{M}$.
To formally obtain the local solution that facilitates labelling the shock of interest to the same column of the impact matrix in all regimes, we recommend employing a ”blended identification” strategy that combines identification by non-Gaussianity with additional identifying information (cf. Carriero et al., 2024). As a concrete example, if a supply shock should increase output and decrease prices at impact in all regimes, selecting the best local solution where the same column in all $B_{1},...,B_{M}$ satisfies such signs amounts to combining identification by non-Gaussianity with the appropriate short-run sign restrictions. In addition to short-run sign restrictions, also other types of information can be exploited in blended identification, including narrative restrictions and testable zero impact effect restrictions, for instance.

### v5

@@ 3.2 Identification of the shocks by non-Gaussianity
    The structural error process $e_{t}=(e_{1t},...,e_{dt})$ is a sequence of independent and identically distributed random vectors such that $e_{t}$ is independent of $\mathcal{F}_{t-1}$ and each component $e_{it}$, i=1,,di=1,...,d, follows a continuous distribution with zero mean, unit variance, and a density that is strictly positive almost everywhere in $\mathbb{R}$.
@@
    The components of $e_{t}=(e_{1t},...,e_{dt})$ are mutually independent and at most one of them has a Gaussian marginal distribution.
The assumption of a unit variance is merely a normalization since the variance of the shocks is captured by the impact matrix $B_{y,t}$. Also, due to the time-variation of the impact matrix, Assumption 1 does not imply that the reduced form innovations $u_{t}$ are independent nor that they are identically distributed. The assumption of strictly positive density on sets of positive (Lebesgue) measure mainly rules out bounded shock distributions.
It is helpful to distinguish two layers of identification. First, for any fixed $t$ and conditional on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ in (3.2) is identified up to column permutations and sign changes by independent component analysis (ICA) under Assumption 1 (cf. Lanne et al., 2017). Second, under the structural parametrization $B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}B_{m}$ in (3.3), identification of the regime-specific matrices $B_{1},...,B_{M}$ (up to column ordering and signs) additionally requires identification of the transition weight parameters, or in the case of exogenous weights, sufficient variation in the known weights. The next lemma, partly similar to Proposition 1 in Lanne et al. (2017), therefore formalizes the per-$t$ ICA result, and, in the case of logistic weights (2.3) establishes identification of the weight function parameters (and, for completeness, identification of the AR parameters for logistic and exogenous weights). Specific weight functions are assumed here for simplicity, but the result can be extended to other suitable weight functions as well. Identification of $B_{1},...,B_{M}$ then follows from these ingredients.
@@
Consider the STVAR model defined in Equations (3.2) and (3.3) with Assumption 1 satisfied. Then, conditionally on $\mathcal{F}_{t-1}$, at each $t$, $B_{y,t}$ is uniquely identified up to ordering and signs of its columns almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
Moreover, suppose the transition weights are either exogenous (nonrandom) or follow the logistic process defined in (2.3) (with $M=2$ assumed), and that the following conditions hold:
@@
    (for exogenous weights) there exists $M$ indices $t_{1},...,t_{M}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, i=1,,Mi=1,...,M, are linearly independent.
Then, the parameters $\phi_{m},A_{m,1},...A_{m,p}$, m=1,,Mm=1,...,M, and for logistic models $c$ and $\gamma$, are uniquely identified.
Lemma 2 is proven in Appendix C.2. Condition (1) implies distinguishable regimes, whereas Condition (2) guarantees a certain small amount of variation in exogenous weights. Lemma 2 concludes that at each $t$, conditionally on $\mathcal{F}_{t-1}$, the impact matrix $B_{y,t}$ is unique up to ordering and signs of its columns under Assumption 1. That is, at each tt‚ changing the ordering or signs of the columns of $B_{y,t}$ would lead to an observationally equivalent model, but changing $B_{y,t}$ in any other way would lead to an observationally distinct model. It follows that if the impact matrix is time-invariant as in Lanne et al. (2017), i.e., $B_{y,t}=B$ for some constant matrix $B$, the structural shocks are identified up to ordering and signs. However, when the impact matrix varies over time, two complications arise.
First, because $B_{y,t}$ (3.3) is identified only up to column ordering and signs at each $t$, its unique identification requires constraints on the regime-specific impact matrices $B_{1},...,B_{M}$ such that any reordering or sign changes in their columns lead to observationally distinct $B_{y,t}$ at some $t$. Second, since $B_{y,t}$ is not a matrix of constant parameters but a function of parameters, it needs to be shown that the parameters in its functional form are identified. Also, in addition to $B_{1},...,B_{M}$, the impact matrix $B_{y,t}$ depends on the transition weights $\alpha_{m,t}$, m=1,,Mm=1,...,M, which must therefore be identified as well. To that end, we assume either the logistic or exogenous transition weights considered in Lemma 2.
The following proposition, proven in Appendix C.3, establishes unique identification of $B_{1},...,B_{M}$.
@@
Consider the STVAR model defined in Equations (3.2) and (3.3) with Assumption 1 and the conditions of Lemma 2 satisfied, and suppose the transition weights are either exogenous (nonrandom) or follow the logistic process defined in (2.3) with $\gamma\in(0,\infty)$ (and $M=2$ assumed). Moreover, suppose the following conditions hold:
@@
    (for exogenous weights) the vector $(\alpha_{1,t},...,\alpha_{M,t})$ takes a value for some $t$ such that none of its entries is zero.
@@
Then, the matrices $B_{1},...,B_{M}$ are uniquely identified almost everywhere in $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$.
@@
Proposition 1 essentially states that if the structural shocks are independent and at most one of them is Gaussian (Assumption 1), the impact matrices $B_{1},...,B_{M}$, and hence, the structural shocks are identified. The result holds for almost every $B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$ because the identification may fail in some special cases, but this set has Lebesgue measure zero. [^2] Beyond possible singularity of the impact matrix $B_{y,t}$ (see Lemma 2), the failure of the identification in a measure zero set only concerns the possibility that fixing the ordering and signs of the columns of $B_{1}$ does not fix the ordering and signs of the columns of $B_{2},...,B_{M}$.
@@
Condition (a) of Proposition 1 states that the ordering and signs of the columns of $B_{1}$ should be fixed (e.g., by assuming that the first nonzero entry in each column is positive and the first nonzero entries are in a decreasing order, given that none of them are equal), which fixes the ordering and signs of the columns of $B_{y,t}$ for all $t$. If the shocks follow, for example, the skewed $t$-distributions defined in Equation (4.1) (in Section 4 discussing estimation), any fixed ordering and signs of the columns of $B_{1}$ can be assumed without loss of generality. This is because reordering the columns would just reorder the shocks and swapping a sign of a column corresponds to swapping the sign of the related shock and skewness parameter. Condition (b) states that exogenous transition weights should, for some $t$, be strictly positive for all the regimes, which is used to establish that Condition (a) fixes the ordering and signs of the columns of $B_{y,t}$ for all $t$. For logistic weights, this is achieved via the assumption $\gamma<\infty$, guaranteeing that the regime-switches are not discrete. [^3] To accommodate discrete regime-switches, which are obtained as a special case of the logistic weights (2.3) with the smoothness parameter tending to infinity, we establish the identification of the shocks in the threshold VAR model of Tsay (1998) in Appendix A.
Somewhat surprisingly, our experience shows that identification appears to be often weak in the sense that there are multiple local maxima of the (penalized) log-likelihood function (discussed in Section 4) with (penalized) log-likelihoods close to each other and to the global maximum. The parameter values related to each such local maximum seem to be often largely quite similar to each other, but have some differences, frequently but not exclusively related to different ordering and signs of the columns of $B_{2},...,B_{M}$. Since such different local solutions may produce different results in structural analysis, this weak identification should be appropriately addressed, for instance, by combining Proposition 1 with supplementary identifying information as discussed in the next section.
3.3 Labeling the shocks and blended identification
As in the linear SVAR model of Lanne et al. (2017), the statistically identified structural shocks do not necessarily have economic interpretations, and labeling them as economic shocks requires external information. However, labeling the shocks based on the estimates of the impact matrices $B_{1},...,B_{M}$ might not always be straightforward, as the same shock must be associated with the same column of the impact matrix $B_{m}$ in all regimes. For example, if a positive supply shock should increase output and decrease prices on impact in all regimes, labeling the $i$th shock as the supply shock requires the $i$th column of all $B_{1},...,B_{M}$ to satisfy such signs. Moreover, as discussed in Section 3.2, our experience shows that identification is often weak in the sense that there are multiple local solutions with fit close to the global optimum. While we find such local solutions to be often largely quite similar to each other, they may produce different impulse response functions even when the shock of interest is associated with the same column of all $B_{1},...,B_{M}$.
To formally address weak identification and the problem of labeling the shocks, we recommend employing a ”blended identification” strategy that combines identification by non-Gaussianity with supplementary identifying information (cf. Carriero et al., 2024). Specifically, we propose imposing overidentifying restrictions that are sufficient to yield a unique local solution (among the ones with fit close to the global solution) and facilitate labeling the shocks of interest. Since different local solutions may yield different results, the restrictions should be economically reasonable and serve to exclude less plausible alternative solutions.
As a simple example, in a bivariate system of output and prices, it may be reasonable to assume that for each variable, one of the shocks has greater impact effect on that variable than the other shock, and that this effect has the same sign across regimes. If one of the shocks is to be interpreted as a demand shock, it might be useful to further assume that the shock with the largest effect on output also moves prices in the same direction in all regimes. Similarly, if the other shock is intended to represent a supply shock, it could be useful to assume that it moves output and prices in the opposite directions in all regimes.444 Also other forms of information can be incorporated into the blended identification strategy, such as narrative restrictions and zero impact effect restrictions (the latter of which are testable due to statistical identification), for instance. See our empirical application in Section 5 for an example that involves more variables.
As a practical consideration, note that our estimation method, described in Section 4 and implemented to the accompanying R package sstvars Virolainen (2025a), produces a set of alternative local solutions. By filtering these local solutions based on whether they satisfy the imposed overidentifying restrictions, it is straightforward to determine which restrictions are sufficient in the considered specification to exclude all but one of the local maximums (among the ones with a fit close to the presumed global optimum). Finally, to label the shocks of interest, the imposed restrictions can often be used to uniquely associate each of them to the same column of the impact matrix across all regimes.